A rocket, initially at rest on the ground, acceler
Answer #1

\large u=0m/s

Initial acceleration

\large a_{in}=53.9m/s^2

\large t=5.0s

\large v=u+at

\large v=53.9m/s^2\times 5sec=269.5m/s until this velocity it will accelerate

\large h_{acc}=\frac{1}{2}a_{in}t^2

\large h_{acc}=673.75

Now it will be the motion under gravity

\large v_f=0

\large (269.5)^2=2\times 9.8\times h_{f}

\large h_{f}=3705.625m

\large y_{max}=H_{total}=h_{acc}+h_f

\large y_{max}=673.75m+3705.625

\large y_{max}=4379.375m